Can you please simplify (9uv)^2(uv)^-3
Answer by dabanfield(803)

Can you please simplify (9uv)^2(uv)^-3
If the problem is:
((9uv)^2)*(uv)^-3 then
We have:
((9^2)*(u^2)*(v^2))/((u^-3)*(v^-3))=
81*u^(2-3)* v(2-3)=
81*(u^-1)*(v^-1) =
81/(u*v)